OpenStudy (anonymous):

please help me solve this problem by using substitution 6x+3y=9 y=−x+4

OpenStudy (anonymous):

6x+3y=9 y=−x+4 multiple equation two in (-3) and then sum it to equation one ok?

OpenStudy (anonymous):

but i still dont get how to do it.. , i wish i was smart, like you guys -.-

OpenStudy (texaschic101):

y = -x + 4 sub -x + 4 in for y in the first equation... 6x + 3(-x + 4) = 9 Now just work it out :)

OpenStudy (texaschic101):

and once you get x...sub it back into either of the original equations and you will get y :)

OpenStudy (anonymous):

so x = 3, and y = -1?

OpenStudy (anonymous):

yup:)